OpenStudy (anonymous):

the average score on a standardized test is 750 points of the standard deviation of 50 points was a probability that a student scores more than 700 standardized test

OpenStudy (anonymous):

\[P(X>700)=P\left(\frac{X-750}{50}>\frac{700-750}{50}\right)=P(Z>-1)\]